OpenStudy (anonymous):

Solve the following equation: -2(x - 1) + 8 ≤ 10 - 4(x + 3)

OpenStudy (linyu):

-2(x-1)+8≤10-4(x+3) -(x-1)+4≤5-2(x+3) -x+1≤1-2x-6 -x+2x≤-6 x≤-6

OpenStudy (texaschic101):

-2(x - 1) + 8 <= 10 - 4(x + 3) -2x + 1 + 8 <= 10 - 4x - 12 -2x + 9 <= -2 - 4x -2x + 4x <= -2 - 9 2x <= - 11 x <= -11/2

OpenStudy (linyu):

-2(x-1) is -2x+2 not -2x+1